A 2-meter copper conductor which has a cross-sectional area of 8 x 10 – 7 m2 has a potential energy difference of 6 volts between its ends. The resistivity of copper is 1.7 x 10 – 8 Ω-m. Find:

a. The resistance of the conductor, and

b. The current passing through it



1
Expert's answer
2021-05-31T02:24:41-0400

"R = \rho\frac{L}{A}"


Where

R = is the resistance

"\rho" = is the resistivity

L= is the length

A = is the cross sectional area


"R= 1.7×10^{-8}\Omega m[\frac{2m}{8×10^{-7}m^2}]"


"R=\frac{3.4×10^{-8}\Omega m^2}{8×10^{-7}m^2}"


"R= 0.0425\Omega"





Current passing


"V= IR"

Where

V = is the potential

I = is the current

R = is the resistance


"I = \frac{V}{R}"


"I =\frac {6V}{0.0425\Omega}= 14.11" amp